OpenStudy (anonymous):

how do graph this equation: (x+3)^2=10(y-1)

OpenStudy (compassionate):

\[ \begin{array}l\color{red}{\text{W}}\color{blue}{\text{E}}\color{red}{\text{L}}\color{blue}{\text{C}}\color{red}{\text{O}}\color{blue}{\text{M}}\color{blue}{\text{E}}\color{red}{\text{ }}\color{blue}{\text{T}}\color{red}{\text{O}}\color{blue}{\text{ }}\color{red}{\text{O}}\color{blue}{\text{P}}\color{blue}{\text{E}}\color{red}{\text{N}}\color{blue}{\text{S}}\color{red}{\text{T}}\color{blue}{\text{U}}\color{red}{\text{D}}\color{blue}{\text{Y}}\color{blue}{\text{!}}\color{red}{\text{ }}\color{blue}{\text{}}\end{array} \]

OpenStudy (compassionate):

FIrst, you want to simplify the equation.

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (compassionate):

Okay, it's easy! \[(x+3)^2=10(y-1)\] Multiple 10 by y and - 1 Then put x to the 2nd power and put 3 to the 2nd power.

OpenStudy (anonymous):

So x^2+9=10y-10?

OpenStudy (compassionate):

Correct! Now tell me a bit about the lesson. Are you studying linear graphs or parabolas?

OpenStudy (anonymous):

Awesome! I am studying "Conic Sections". That is the name of the lesson.

OpenStudy (compassionate):

Okay, the equation graphed looks something like this. (Forgive my bad drawing. Lol) 

OpenStudy (compassionate):

Here: http://prntscr.com/10geln

OpenStudy (anonymous):

So would the line of symmetry be the x axis? (the question asks for the line of symmetry)

OpenStudy (compassionate):

The line of syemtry is the axis, yes!

OpenStudy (anonymous):

Awesome! Thanks!